Find the point of the line y=x nearest to the point (4,1)


1
Expert's answer
2021-01-19T02:19:30-0500

Any point in the line y=x is (x,x), where x is a real number.

Distance between (x,x) and (4,1) equals:


"d(x)=\sqrt{(x-4)^2+(x-1)^2}"

"d(x)=\sqrt{x^2-8x+16+x^2-2x+1}"

"d(x)=\sqrt{2x^2-10x+17}"


We need to find the minimum of d(x). Square root function is increasing, so finding the minimum of d(x) is equivalent to finding the minimum of square root's argument.


"f(x)=2x^2-10x+17"


Find the extremum point:


"f'(x)=0"

"4x-10=0"

"x=2.5"


To determine if this point is a minimum we calculate the second derivative at this point.


"f''(x)=4"

"f''(2.5)=4>0"


The second derivative is greater than zero, so x=2.5 is minimum point of f(x) and d(x).

So, point (2.5, 2.5) is the nearest point in the line y=x to the point (4, 1).